OpenStudy (anonymous):

The median of the number set below is 40. 15 , 34 , 53 , 62 , 58 , 40 , 21 If the number 45 is added to the data set, what is the new median? 46.5 45 42.5 40

OpenStudy (anonymous):

@skyelaaboss

OpenStudy (bibby):

order the numbers from least to greatest if there's one number in the middle, that's your median if there are 2, you take their average for example the median in 1, 2, 4, 6, 10 is 4 but the median in 10 15 20 30 45 55 is 20+30/2=25

OpenStudy (anonymous):

A?

OpenStudy (anonymous):

The median is 42.5

OpenStudy (bibby):

40+45/2=85/2=42.5

OpenStudy (anonymous):

You put them all in order including the 45 and you have two middle numbers (40,45) then you add them together and divide by 2 and you get 42.5

OpenStudy (anonymous):

thank you both!

OpenStudy (bibby):

15, 21 , 34 , 40 , 53 , 58 , 62 add 45 \(\color{red}{15}, \color{red}{21} , \color{red}{34}, 40 , 45 , \color{red}{53}, \color{red}{58}, \color{red}{62}\)

OpenStudy (anonymous):

No problem(:

OpenStudy (bibby):

npnp